OpenStudy (anonymous):

help factoring!! the problem is -16x2 + 60x + 16 then i factored the GCF of -4 and got -4(4x^2 - 15x - 4) but i can't seem to find nmbers that multiply to -4 and add to -15. am i right? or is multiply to 15 and add to 4? i need to find the zeros.

OpenStudy (anonymous):

@math&ing001

OpenStudy (anonymous):

@jim_thompson5910

myininaya (myininaya):

multiply to be -16 and add up to be -15 is what you mean?

OpenStudy (anonymous):

yea, i need to split the middle term

myininaya (myininaya):

well what is -16(1) and -16+1 equal to?

OpenStudy (anonymous):

original: -16x2 + 60x + 16 my gess after factoring GCF: -4(4x^2 - 15x - 4)

OpenStudy (anonymous):

no, is -15 and -4

myininaya (myininaya):

no...You are to find two factors of a*c that have product ac and sum b.

myininaya (myininaya):

you need to find two number that multiply to be -16 and add up to be -15

OpenStudy (anonymous):

ohh.....

OpenStudy (anonymous):

thanks, i was in the wrong path lol

myininaya (myininaya):

not entirely you were able to factor out that -4 from the start :)

OpenStudy (anonymous):

lol true, well thank you!!

myininaya (myininaya):

let me know if you have trouble

OpenStudy (anonymous):

okay i get: -4(4x+1)(x-4) correct me if im wrong, im supposed to find the zeros... 4, -1/4, and 4 ?

OpenStudy (anonymous):

@myininaya

myininaya (myininaya):

looks great! :)

myininaya (myininaya):

and the zeros...why do you have three of them

OpenStudy (anonymous):

because -4 is factored out the parenthesis and i equaled to it zero

myininaya (myininaya):

one is just a repetition but the x=4 only occurs once so you have the zeros are x=-1/4 or x=4

myininaya (myininaya):

-4 can never be 0

myininaya (myininaya):

4x+1 can be zero x-4 can be zero but -4 can never be zero

OpenStudy (anonymous):

oh okay i see

myininaya (myininaya):

i still think you did pretty well so i will give you a medal for your hard and good work

OpenStudy (anonymous):

one more, correct me if im wrong, since the leading is negative, the parabola is going to open.... down?

OpenStudy (anonymous):

thanks

myininaya (myininaya):

very good

OpenStudy (anonymous):

just making sure the info was frozen in my head, had to take this exam a few days ago so needed a refresher

myininaya (myininaya):

 it has two x-intercepts x=4 and x=-1/4 as the graph shows and it is open down because the number in front of the x^2 is neg

OpenStudy (anonymous):

i see

OpenStudy (anonymous):

yea the zeros are x-intercepts